What is the final pressure, in atmospheres, when the fire extinguisher is used at a temperature of 81 C

, if volume and amount of gas do not change?


1
Expert's answer
2021-10-24T01:32:52-0400

A fire extinguisher has a pressure of 12 atm at 25 ∘C. What is the final pressure, in atmospheres, when the fire extinguisher is used at a temperature of 81 ∘C , if volume and amount of gas do not change?

PV = nRT

P is directly proportional to temperature .

"P_1 = 12 \; atm \\

0°C = 273.15 \;K \\

T_1 = 25°C = 25+273.15 \;K = 298.15 \;K \\

T2 = 81°C = 81+273.15 \;K = 354.15 \; K \\

\frac{P_1}{T_1} = \frac{P_2}{T_2} \\

\frac{12}{298.15} = \frac{P_2}{354.15} \\

0.04024 = \frac{P2}{354.15} \\

P_2 = 0.04024 \times 354.15 = 14.25 \; atm"